OpenStudy (anonymous):

Ratio: there are 3 yes votes to every 4 no votes, how many total votes are there if 2754 people vote yes?

OpenStudy (anonymous):

soo, im not positive on how to do this but i'll give it a shot

myininaya (myininaya):

y/n=3/4 y+n=2754

OpenStudy (anonymous):

Give it a shot, I have 6425

myininaya (myininaya):

oops wait y=2754 sorry y/n=3/4

myininaya (myininaya):

y+n will be ur answer

OpenStudy (anonymous):

sorry 6426

OpenStudy (anonymous):

soo, maybe (4*2754)/3 = n

OpenStudy (anonymous):

But what is n?

OpenStudy (anonymous):

then add that to 2754

myininaya (myininaya):

n is number of no votes

OpenStudy (anonymous):

no, i mean what value do i plug in for n?

myininaya (myininaya):

thats the unknown you know what y is solve for n

OpenStudy (anonymous):

you need to solve for n first in that equation that i gave

OpenStudy (anonymous):

i also got 6426. so you're right

OpenStudy (anonymous):

i hope, i divided 2754 by 3, then multiplied by 4

OpenStudy (anonymous):

yup. then added 2754 again.

myininaya (myininaya):

n=3672 y+n=2754+3672=6426